The pressure reduction of a sample of 27 fuel valves in a preliminary test sample for potential use in heart bypass surgeries showed a standard deviation of 0.04 ounces. The manufacturer claims the population variance is less than 0.002.( Ho: σ2 The test statistic is (Round your answer to 3 decimal places.) 0.002) Answer 1 Point Keypad Tables !
Answer #1

Solution:- Given that information: standard deviation = s = 0.04 , ?^2 = 0.002 n = 27

Test statistic = X^2 = (n-1)*(s^2/?^2)
= (27-1)*(0.04^2/0.002)
= 20.800